OpenStudy (anonymous):

Find the equation of a line that goes through points (0, 4) and (1, 6)

OpenStudy (anonymous):

You want to find the equation for a line that passes through the two points: (0,4) and (1,6). Use y=mx+b First: find the slope y2-y1 6-4 2 ----- = ---- = -- x2-x1 1-0 1 m=2 therefor y=2x+4